Calculate the probabilities of P(-z0<z<0)= 0.3708
1
Expert's answer
2021-09-30T23:47:29-0400

"P(-z_0<Z<0)=0.5-P(Z<-z_0)"

"0.5-P(Z<-z_0)=0.3708"

"P(Z<-z_0)=0.5-0.3708=0.1292"

"-z_0=-1.13"

"z_0=1.13"